Keith bought 30 lb of nuts at $3.00/lb and 20 lb of nuts at $5.00/lb. If he wanted to buy 10 more pounds of a different kind of nut to make the average price per pound equal to $4.50, what price should he pay for the additional 10 lb?
Please help :)

Answer by CubeyThePenguin(3113)

(30 * $3 + 20 * $5)/50 = $3.80 (average price of mixture at the start)

(50 * $3.80 + 10x)/60 = $4.50
$190 + 10x = $270
10x = 80
x = 8

He needs to pay for nuts at $8/lb.